# In-plane Black-hole Spin Measurements Suggest Most Gravitational-wave Mergers Form in Triples

### v1

@@ Acknowledgments
We thank Colm Talbot, Isobel Romero-Shaw, and Carlos Lousto for useful discussions and input. S.R. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 546677095. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@
Appendix A Supplementary results about the Gaussian + Isotropic + Cut model
In Figure 4 we show a corner plot of selected parameters of the Gaussian + Isotropic + Cut model, showing that the near-perpendicular Gaussian component ($\mu_{t}=0.20^{+0.21}_{-0.11}$ and $\sigma_{t}=0.55^{+0.25}_{-0.16}$) is the dominant component ($\xi=0.86^{+0.10}_{-0.55}$) at low masses ($m_{1}\lesssim\tilde{m}=44.26^{+8.65}_{-4.57}\,\rm M_{\odot}$).
Figure 4: Joint and marginal posteriors for the mean $\mu_{t}$ and standard deviation $\sigma_{t}$ of the Gaussian component, its mixing fraction $\xi$, and the mass cut-off $\tilde{m}$ in the Gaussian + Isotropic + Cut model.
Appendix B Other population models
@@
Figure 5 shows the marginal posterior distributions of the Gaussian mixing fraction $\xi$. In all models, where the Gaussian is allowed to move freely ($1\leq\mu_{t}\leq 1$) it dominates the distribution, while allowing for a potential mixing with a subdominant isotropic component. Enforcing a preferentially aligned spin-orbit configuration ($\mu_{t}=1$ in Aligned + Isotropic and Aligned + Isotropic + Cut) makes the Gaussian component a subdominant contribution, suggesting no preference for alignment in the gravitational-wave data. Figure 5 shows only small probabilities for the aligned components to be dominant, e.g., the cumulative density function yields $P(\xi\lesssim 0.8)\approx 0.95$ and $0.85$ in the Aligned + Isotropic) and Aligned + Isotropic + Cut models, respectively. However, Figure 6 shows that a large aligned contribution $\xi\lesssim 1$ is correlated with large width $\sigma_{t}\approx 1$ (which is different to the Gaussian + Isotropic + Cut model in Figure 4). Thus, we conclude that the Bayes factors disfavour aligned models at varying degree (see Section 3) and even if realised they are either a subdominant component (Figure 5) or they dominate but are too wide to be plausibly explained by isolated binary formation scenarios (see Section 4) and indistinguishable from the isotropic component (see Figure 8 below).
@@
In Figure 7, we further investigate the preference against an aligned component. We explore variants where the aligned component is restricted to $\cos\theta\in[t_{\min},1]$ for some fixed threshold value $t_{\min}$, since realistic isolated binary evolution scenarios may yield spin-orbit tilts which are more narrowly confined than the entire range $\cos\theta\in[-1,1]$. For this purpose, we adopt the Aligned + Isotropic + Cut model but use truncated normal distributions $\mathcal{N}_{[t_{\min},1]}(\cos\theta_{i}|\mu_{t}=1,\sigma_{t})$ in Equation (2) with $t_{\min}=-1.0,-0.9,-0.8,\dots,0.9$. The left panel of Figure 7 shows that larger values of $t_{\min}$ (i.e., more narrowly distributed spin-orbit angles about alignment) require smaller contributions from the aligned component, ranging from $\xi(t_{\min}=-1)=0.09^{+0.66}_{-0.06}$ to $\xi(t_{\min}=0.9)=0.01\pm 0.01$. In the right panel, we compare the modified Aligned + Isotropic + Cut models to the Gaussian + Isotropic + Cut model (where the Gaussian component remains confined to $\cos\theta\in[-1,1]$). Evidence against the aligned models is positive (Kass) with $|\Delta\ln\mathcal{B}|\approx 1$ to 3 across the whole range of $t_{\min}$. We also might see a tendency that above $t_{\min}\gtrsim 0.3$ models become more strongly disfavoured, which would also roughly coincide with the depletion of spin-orbit tilt angles in the non-parametric B-Spline model (cf. Figure 2). However, we do note that the scatter in $\Delta\ln\mathcal{B}$ across different explored models is rather large and inhibits definite conclusions. Since $\Delta\ln\mathcal{B}$ scales roughly with the detected sample size (Kass) we expect the robustness of the model comparison to improve soon as the gravitational-wave catalogue grows during the fourth and fifth LVK observing runs.
@@
Figure 7: Aligned + Isotropic + Cut model for different lower bounds on $\cos\theta_{i}$ of the aligned component, i.e., $\mathcal{N}_{[t_{\min},1]}(\cos\theta_{i}|\mu_{t}=1,\sigma_{t})$ in Equation (2) with $t_{\min}=-1.0,-0.9,-0.8,\dots,0.9$. The left panel shows the marginal posterior distribution of the contribution $\xi$ of the low-mass aligned component (whereas $1-\xi$ corresponds to the low-mass isotropic component). The right panel shows the strength of evidence against each aligned model compared to our default Gaussian + Isotropic + Cut model, which allows but does not enforce preferentially aligned mergers. Nomenclature follows Kass.
In addition, Figure 8 shows that that the PPD of the spin-orbit tilts in most models with a free Gaussian recover the pronounced peak at $\cos\theta\approx 0$ of the non-parametric model (GWTC4pop). The peak is less pronounced and recovered in models that assume a uniform prior on $\sigma_{t}$ (Uniform Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic), where shallow distributions with large values of $\sigma_{t}$ mimic an isotropic distribution. As discussed in Section 2, a uniform prior on $\sigma_{t}$ is not expected from astrophysical source modelling. We note that LVK Uniform Gaussian + Isotropic is a similar to the Gaussian + Isotropic model by GWTC4pop and recovers a similar PPD.

### v2

@@ Acknowledgments
We thank the two anonymous reviewers, Colm Talbot, Isobel Romero-Shaw, and Carlos Lousto for useful comments and input that helped improving this work. FA and VR are supported by the UK’s Science and Technology Facilities Council grants ST/V005618/1 and UKRI2489. S.R. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 546677095. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
Appendix A Mass and redshift modelling
In all our models, we assume a simple parameterisation of our mass distribution as (Talbot et al., 2019, 2025)
$\pi(m_{1},q)=\pi(m_{1})\pi(q|m_{1}),$ (A1)
where $0<q\leq 1$ is the binary mass ratio and
$\pi(m_{1})$ $=(1-\lambda_{m})\pi_{\text{pow}}(m_{1})+\lambda_{m}\pi_{\text{norm}}(m_{1}),$ (A2)
$\pi_{\text{pow}}(m_{1})$ $\propto m^{-\alpha}\quad\text{with}\quad m_{\min}\leq m_{1}<m_{\max},$ (A3)
$\pi_{\text{norm}}(m_{1})$ $\propto\exp\left(-\frac{(m_{1}-\mu_{m})^{2}}{2\sigma^{2}_{m}}\right)$
$\quad\text{with}\quad m_{\min}\leq m_{1}<100\,\rm M_{\odot},$ (A4)
$\pi(q|m_{1})$ $=\frac{1+\beta}{1-(m_{\min}/m_{1})^{1+\beta}}q^{\beta}.$ (A5)
Thus, the primary mass distribution is modelled as a power-law with a Gaussian component and the conditional mass-ratio distribution as a power-law.
For the distribution in redshift $z$ we assume a power-law following Fishbach et al. (2018)
$\pi(z)=(1+z)^{\lambda}.$ (A6)
For all hyper-parameters introduced in Equations (A2) – (A6) we assume uniform priors within the bounds given in Table 2 (Talbot et al., 2019, 2025).
Parameter Description Prior
$\lambda_{m}$ $m_{1}$ mixture fraction $\mathcal{U}_{[0,1]}$
$\alpha$ $m_{1}$ power-law index $\mathcal{U}_{[-2,4]}$
$m_{\min}/\mathrm{M}_{\odot}$ minimum mass $\mathcal{U}_{[2.0,2.5]}$
$m_{\max}/\mathrm{M}_{\odot}$ maximum mass $\mathcal{U}_{[80,100]}$
$\mu_{m}/\mathrm{M}_{\odot}$ $m_{1}$ Gaussian component peak $\mathcal{U}_{[10,50]}$
$\sigma_{m}/\mathrm{M}_{\odot}$ $m_{1}$ Gaussian component width $\mathcal{U}_{[1,10]}$
$\beta$ mass ratio power-law index $\mathcal{U}_{[-4,12]}$
$\lambda$ redshift power-law index $\mathcal{U}_{[-1,10]}$
Table 2: Prior distributions for mass and redshift model hyper-parameters.
We emphasise that for models that mix two components (3 – 9 in Table 1) we assume the same mass and redshift modelling for both, which is consistent with the approach of the LVK collaborations (e.g., The LIGO, Virgo, and KAGRA Collaborations, 2023b, 2025c). However, if both components represent two different astrophysical formation channels (e.g., a near-perpendicular Gaussian from triple star evolution and an isotropic component from dynamical interactions in star clusters) they could in reality follow different mass and redshift distributions (cf., Dorozsmai et al., 2025). Investigating the impact of our simplification as well as allowing for any spin correlation with mass or redshift are left for future work.
Appendix B Supplementary results about the Gaussian + Isotropic + Cut model
@@
Figure 4: All joint and marginal posteriors in the Gaussian + Isotropic + Cut model. Variables are defined in Section II and Appendix A.
Figure 5: PPD of the cut-off Equation (4) in the Gaussian + Isotropic + Cut model. Blue shows the PPD for posterior draws of $\tilde{m}$, and orange shows the $90\%$ bounds for prior draws of $\tilde{m}$ (uniform distribution $\mathcal{U}_{[10\,\rm M_{\odot},100\,M\odot]}$, see Section II).
Figure 6: Same as Figure 4 with a free parameter $\tilde{k}$.
Appendix C Other population models
@@
Figure 7 shows the marginal posterior distributions of the Gaussian mixing fraction $\xi$. In all models, where the Gaussian is allowed to move freely ($1\leq\mu_{t}\leq 1$) it dominates the distribution, while allowing for a potential mixing with a subdominant isotropic component. Enforcing a preferentially aligned spin-orbit configuration ($\mu_{t}=1$ in Aligned + Isotropic and Aligned + Isotropic + Cut) makes the Gaussian component a subdominant contribution, suggesting no preference for alignment in the gravitational-wave data. Figure 7 shows only small probabilities for the aligned components to be dominant, e.g., the cumulative density function yields $P(\xi\lesssim 0.8)\approx 0.95$ and $0.85$ in the Aligned + Isotropic) and Aligned + Isotropic + Cut models, respectively. However, Figure 8 shows that a large aligned contribution $\xi\lesssim 1$ is correlated with large width $\sigma_{t}\approx 1$ (which is different to the Gaussian + Isotropic + Cut model in Figure 4). Thus, we conclude that the Bayes factors disfavour aligned models at varying degree (see Section III) and even if realised they are either a subdominant component (Figure 7) or they dominate but are too wide to be plausibly explained by isolated binary formation scenarios (see Section IV) and indistinguishable from the isotropic component (see Figure 10 below).
@@
In Figure 9, we further investigate the preference against an aligned component. We explore variants where the aligned component is restricted to $\cos\theta\in[t_{\min},1]$ for some fixed threshold value $t_{\min}$, since realistic isolated binary evolution scenarios may yield spin-orbit tilts which are more narrowly confined than the entire range $\cos\theta\in[-1,1]$. For this purpose, we follow a similar approach to Callister et al. (2022); Tong et al. (2022) and adopt the Aligned + Isotropic + Cut model but use truncated normal distributions $\mathcal{N}_{[t_{\min},1]}(\cos\theta_{i}|\mu_{t}=1,\sigma_{t})$ in Equation (2) with $t_{\min}=-1.0,-0.9,-0.8,\dots,0.9$. The left panel of Figure 9 shows that larger values of $t_{\min}$ (i.e., more narrowly distributed spin-orbit angles about alignment) require smaller contributions from the aligned component, ranging from $\xi(t_{\min}=-1)=0.09^{+0.66}_{-0.06}$ to $\xi(t_{\min}=0.9)=0.01\pm 0.01$. In the right panel, we compare the modified Aligned + Isotropic + Cut models to the Gaussian + Isotropic + Cut model (where the Gaussian component remains confined to $\cos\theta\in[-1,1]$). Evidence against the aligned models is positive (Kass and Raftery, 1995) with $|\Delta\ln\mathcal{B}|\approx 1$ to 3 across the whole range of $t_{\min}$. We also might see a tendency that above $t_{\min}\gtrsim 0.3$ models become more strongly disfavoured, which would also roughly coincide with the depletion of spin-orbit tilt angles in the non-parametric B-Spline model (cf. Figure 2). However, we do note that the scatter in $\Delta\ln\mathcal{B}$ across different explored models is rather large and inhibits definite conclusions. Since $\Delta\ln\mathcal{B}$ scales roughly with the detected sample size (Kass and Raftery, 1995) we expect the robustness of the model comparison to improve soon as the gravitational-wave catalogue grows during the the fourth and fifth, as well as a possible intermediate, LVK observing runs.
@@
Figure 9: Aligned + Isotropic + Cut model for different lower bounds on $\cos\theta_{i}$ of the aligned component, i.e., $\mathcal{N}_{[t_{\min},1]}(\cos\theta_{i}|\mu_{t}=1,\sigma_{t})$ in Equation (2) with $t_{\min}=-1.0,-0.9,-0.8,\dots,0.9$. The left panel shows the marginal posterior distribution of the contribution $\xi$ of the low-mass aligned component (whereas $1-\xi$ corresponds to the low-mass isotropic component). The right panel shows the strength of evidence against each aligned model compared to our default Gaussian + Isotropic + Cut model, which allows but does not enforce preferentially aligned mergers. Nomenclature follows Kass and Raftery (1995).
In addition, Figure 10 shows that that the PPD of the spin-orbit tilts in most models with a free Gaussian recover the pronounced peak at $\cos\theta\approx 0$ of the non-parametric model (The LIGO, Virgo, and KAGRA Collaborations, 2025c). The peak is less pronounced and recovered in models that assume a uniform prior on $\sigma_{t}$ (Uniform Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic), where shallow distributions with large values of $\sigma_{t}$ mimic an isotropic distribution. As discussed in Section II, a uniform prior on $\sigma_{t}$ is not expected from astrophysical source modelling. We note that LVK Uniform Gaussian + Isotropic is a similar to the Gaussian + Isotropic model by The LIGO, Virgo, and KAGRA Collaborations (2025c) and recovers a similar PPD.